A flower pot from a roof top garden falls off the edge of the roof and falls past the window below.Air resistance may be ignored.The flower pot takes 0.48seconds to travel between the topand the bottom of the window which is 1.9meters high.How far below the roof is the top of the window?
1
Expert's answer
2020-02-24T10:50:04-0500
"h=1.9=h_2-h_1=0.5gt_2^2-0.5gt_1^2=0.5g(t_2^2-t_1^2)"

"2h=g(t_2-t_1)(t_2+t_1)=g(t)(t_2+t_1)"

"t_2-t_1=t=0.48\ s"

We need to find "h_1".


"2(1.9)=9.8(0.48)(t_2+t_1)"

"(t_2+t_1)=0.808\ s"

"t_1=0.5((t_2+t_1)-t)=0.5(0.808-0.48)=0.164\ s"

"h_1=0.5(9.8)(0.164)^2=0.13\ m"